Can someone show me how to do this one?
Business and Finance. The demand equation for a certain type of printer is given by
D= -200p + 35,000
The supply equation is predicted to be
S= -p^2 + 400p - 20,000
Find the eqilibrium price.
Thank you!

Answer by scott8148(6628)

equilibrium price is when supply equals demand

-p^2+400p-20000=-200p+35000 ... 0=p^2-600p+55000